OpenStudy (anonymous):

the population of humorville is 8,500 people. in this town, jokes travel fast. in one hour, each person who hears a joke tells three other people who have not heard it, and tells no one else. last frieday, a visitor from out of town told the police chief a new joke at 10:00 a.m. how long did it take for everyone in humorville to hear the joke?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

so, what are your thoughts?

OpenStudy (anonymous):

my thoughts is that this is an exponent problem @amistre64

OpenStudy (amistre64):

it is an exponent yes have you tried making a list of values? h0 = 1 h1 = 1 + 3 h2 = 1 + 3 + 3(3) h3 = 1 + 3 + 3(3) + 3(3)(3) etc ....

OpenStudy (mathstudent55):



OpenStudy (anonymous):

no i havent @amistre64

OpenStudy (amistre64):

well, making a list helps to see the process

OpenStudy (amistre64):

what does my list suggest to you?

OpenStudy (mathstudent55):

This is a geometric sequence which is exponential, but it's more than that. In other words, you need to find after which period, the sum of the sequence reaches 10,000.

OpenStudy (anonymous):

to add and multiply @amistre64

OpenStudy (amistre64):

doesnt the number of people for a given hour look like a geometric sum? if so, how do we solve a geometric summation?

OpenStudy (amistre64):

\[h_n=\sum_{k=0}^{n}3^n\] when does this equate to the entire population?

OpenStudy (amistre64):

3^k that is

OpenStudy (anonymous):

wait im confused @amistre64

OpenStudy (amistre64):

i hope you get more specific than that

OpenStudy (anonymous):

i dont understand how to solve a geometric summation

OpenStudy (amistre64):

well, with any luck you have a formula laying about for it

OpenStudy (amistre64):

the other method is to simply do something nifty

OpenStudy (amistre64):

S = 1 + r + r^2 + r^3 + r^4 + ... + r^n -rS = - r - r^2 - r^3 - r^4 + ... + r^n - r^(n+1) when we add these together, almost everything goes to 0

OpenStudy (amistre64):

S = 1 + r + r^2 + r^3 + r^4 + ... + r^n -rS = - r - r^2 - r^3 - r^4 + ... + r^n - r^(n+1) --------------------------------------------- (1-r)S= 1 + 0 + 0 + 0 + 0 + ... + 0 -r^(n+1) out goal is to find S so, divide both sides by 1-r \[S=\frac{1-r^{n+1}}{1-r}\]

OpenStudy (amistre64):

so now we have a formula to determine the summation of any nth hour. how do we determine the solution to the problem now?